OpenStudy (anonymous):

Find the slope of the line that passes through the points (4, 8) and (6, -2).

OpenStudy (amistre64):

repeat lol

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) = (-2-8)/(6-4) =-10/2 =-5